OpenStudy (anonymous):

trig id's verify:

OpenStudy (anonymous):

\[\frac{ \cot(x)\sec(x) }{ \csc(x) } = 1\]

OpenStudy (anonymous):

\[\frac{ (\cos(x)/\sin(x))(1/\cos(x) }{ 1/\sin(x) }\]

OpenStudy (anonymous):

\[(\frac{ \cos(x)/\sin(x) }{ 1/\sin(x) })(\frac{ 1/\cos(x) }{ 1/\sin(x) })\]

OpenStudy (anonymous):

\[(\frac{ \cos(x) }{ \sin(x) } \times \frac{ \sin(x) }{ 1 })(\frac{ 1 }{ \cos(x) }\times \frac{ \sin(x) }{ 1 })\]

OpenStudy (anonymous):

\[(\cos(x))(\cot(x))\]

OpenStudy (anonymous):

does not equal 1 right

OpenStudy (anonymous):

Not sure where you got the cot(x) from, but you are welcome to show me just in case i'm wrong. \[\frac{ \frac{ \cos(x) }{ (\sin(x) } }{ \frac{ 1 }{ \sin(x) } } = \frac{ \cos(x) }{ 1 } = \cos(x)\]\[\frac{ \frac{ 1 }{ \cos(x) } }{ \frac{ 1 }{ \sin(x) } } = \frac{ \frac{ 1 }{ \sin(x)\cos(x) } }{ \frac{ 1 }{ \sin(x)\cos(x) } } = 1\]\[\cos(x) \times 1 = \cos(x) \neq 1\]

OpenStudy (anonymous):

oh my bad i thot cot x = sin x/cos x when it should be cot x = cos x/sin x

OpenStudy (anonymous):

thanks for your time sir^^)

OpenStudy (anonymous):

Let me check the whole thing again. I am second guessing myself.

OpenStudy (anonymous):

no you are right! i made a simple blind mistake i need glasses!